OpenStudy (anonymous):

PLEASE HELP WILL FAN AND GIVE MEDAL the sum of two consecutive odd integers is 244. what is the smaller integer. A. 119 B. 125 C. 123 D. 121

OpenStudy (jonnyvonny):

They tell you the smaller integer, or whole number, by giving you the smaller one as answer choices. 119 + 121 = 125 + 127 = 123 + 125 = 121 + 123 = From these, which equals 244? Post your response, and I'll post mine

OpenStudy (anonymous):

3(x-1) - (x+5) = 2 ; distribute the 3 and the - sign 3x-3 - x - 5 = 2; combine like terms 2x - 8 = 2 ; add 8 to each side 2x=8+2; do some addition 2x=10; divide each side by 2 x=5; answer (A) 121+123 = 244; so D 3. Make an equation: Let x = daughters age 36 = 3x-6; add 6 to both side 42 = 3x; divide by 3; 14=x Answer C

OpenStudy (jonnyvonny):

@crazyjay100 , so, you do realize that you put in more effort than him in solving his question... right?

OpenStudy (anonymous):

thanks